# Optimal Bound for PCA with Outliers using Higher-Degree Voronoi Diagrams

### v2

@@ Abstract
In this paper, we introduce new algorithms for Principal Component Analysis (PCA) with outliers. Utilizing techniques from computational geometry, specifically higher-degree Voronoi diagrams, we navigate to the optimal subspace for PCA even in the presence of outliers. This approach achieves an optimal solution with a time complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$. Additionally, we present a randomized algorithm with a complexity of $2^{\mathcal{O}(r(d-r))}\times\text{poly}(n,d)$. This algorithm samples subspaces characterized in terms of a Grassmannian manifold. By employing such sampling method, we ensure a high likelihood of capturing the optimal subspace, with the success probability $(1-\delta)^{T}$. Where $\delta$ represents the probability that a sampled subspace does not contain the optimal solution, and $T$ is the number of subspaces sampled, proportional to $2^{r(d-r)}$. Our use of higher-degree Voronoi diagrams and Grassmannian based sampling offers a clearer conceptual pathway and practical advantages, particularly in handling large datasets or higher-dimensional settings.
@@ 1 Introduction
$\begin{split}&\texttt{Input:}\>\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R}),r\in\{1,2,\cdots,d\}\\ &\arg\min_{L}||\text{X}-L||_{F}^{2}\ \\ &\texttt{Subject}\>\texttt{to:}\>rank(L)\leq r\end{split}$ (1)
Here $||\text{A}||_{F}^{2}=\sum_{i,j}{a_{ij}^{2}}$ is the square of the Frobenius norm of matrix A.
@@
In the presence of outliers, the classical PCA algorithm may not work very well, as they can significantly affect the results. The PCA with Outliers problem attempts to improve the results of PCA by removing some data points that are considered outliers. The goal is to recover a low-rank matrix with sparse errors. Here, the assumption is that the data points lie near an $r$-dimensional subspace, and the remaining points are distributed randomly, possibly as a result of distorted observations. This problem can be formulated as follows: given a matrix $\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R})$, where $n$ is the number of data points and $d$ is the dimensionality of the data, and a parameter $r$ representing the rank of the low-rank matrix, the goal is to find a low-rank matrix $L$ and a sparse error matrix $S$ such that $\text{X}=S+L$, where $rank(L)\leq d-r$ and $||S||_{0}\leq k$, where $k$ is a parameter representing the number of outliers to be removed, and $||\cdot||_{0}$ denotes the number of zero entries in a matrix, as shown in (2).
$\begin{split}&\texttt{Input:}\>\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R}),r\in\{1,2,\cdots,d\},\\ &k\in\{1,2,\cdots,n\}\\ &\texttt{Output:}\arg\min_{S,L}||\text{X}-S-L||_{F}^{2}\\ &\texttt{Subject}\>\texttt{to:}\>rank(L)\leq r,||S||_{0}\leq\mathtt{k}\end{split}$ (2)
@@
Our Results. The contribution of this paper is as follows. First, we propose a new algorithm for PCA with Outliers with time complexity $n^{\mathcal{O}(d)}\text{poly}(n,d)$, which fills the gap identified in [26] as this is the optimal bound. To achieve this, we leverage higher-degree Voronoi diagrams, which are defined in the next section. Second, we develop a randomized algorithm with time complexity $2^{O(r(d-r))}\text{poly}(n,d)$. This algorithm provides an optimal solution for PCA with Outliers with a probability of at least $(1-\delta)^{T}$, where $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$ and $T\propto 2^{r(d-r)}$. Here, $\alpha$ is a constant parameter, and $T$ is the number of subspaces sampled. Also, $\delta$ is defined as the probability of an open ball centered around point $L$ with radius $\alpha$ in the Grassmannian manifold, all to be defined in the later sections. As $\alpha$ decreases, $T$ increases exponentially, ensuring a high success probability. This approach offers a more intuitive solution compared to [26], incorporating an effective sampling technique to accurately hit the optimal subspace. Specifically, we employ a probabilistic method to sample subspaces from the Grassmannian manifold, ensuring that with high probability, the sampled subspaces are close to the optimal solution.
@@ 1.1 Related works
Where $\|\cdot\|_{*}$ denotes the nuclear norm of a matrix, which is the sum of its singular values, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm of a matrix, which is the sum of the absolute values of its entries. The parameter $\lambda$ controls the trade-off between the low-rank and sparse components of the data matrix. The PCA with Outliers problem can be solved using a variety of convex optimization techniques, such as the Alternating Direction Method of Multipliers or the Augmented Lagrangian Method.
The paper [29] proposes the Outlier Pursuit method to solve this convex problem. This algorithm is based on the observation that the sparse matrix $S$ can be decomposed into a sum of sparse vectors, each of which corresponds to an outlier. It iteratively identifies and removes the outliers from the data matrix, and then computes the low-rank component of the remaining data using PCA. The outliers are then added back to the low-rank component to obtain the final estimate of the data matrix. The mentioned algorithm requires the sparsity level to be known a priori, which can be a limitation in practice. Additionally, the algorithm assumes that the noise in the data matrix is Gaussian, which may not be true in practice.
To overcome the shortcomings of the previous methods, [9] proposed the Manipulator Pursuit algorithm, which is more robust and flexible. Manipulator Pursuit can handle a wider range of outlier models and noise distributions. It does not require the sparsity level to be known a priori as it uses a data-driven approach to estimate the sparsity level and a nuclear norm regularization term to encourage low-rank solutions.
In [26], the authors analyze the computational complexity of the PCA with Outliers problem. The authors refine the lower bound on the time complexity of the problem, showing that it is $\Omega(n^{d})$. The authors conclude that the lower bound is tight up to a polynomial factor, and also propose an algorithm solving this problem in $n^{\mathcal{O}(d^{2})}{\text{poly}(n,d)}$ using tools from real algebraic geometry. To be explicit, the result is summarized in the following proposition.
@@
([5], Theorem 13.22) Consider $V$ as an algebraic set in $\mathbb{R}^{d}$ of real dimension $d^{\prime}$, defined by $Q(X_{1},\ldots,X_{d})=0$, where $Q$ is a polynomial in $\mathbb{R}[X_{1},\ldots,X_{d}]$ of degree at most $b$. Let $\mathcal{P}\subset\mathbb{R}[X_{1},\ldots,X_{d}]$ be a finite set of $s$ polynomials, each $P\in\mathcal{P}$ also having a degree at most $b$. Define $D$ as the ring generated by the coefficients of $Q$ and the polynomials in $\mathcal{P}$. An algorithm exists that, given $Q$, $d^{\prime}$, and $\mathcal{P}$, computes a set of points satisfying every non-empty cell of $V$ over $\mathcal{P}$. This algorithm requires at most $s^{d^{\prime}}b^{O(d)}$ arithmetic operations in $D$.
@@
    For each $V_{C}$, compute the optimal $S_{V_{C}}$ by selecting the $k$ rows of $X$ with the highest value of $||V_{C}x^{T}_{i}||_{F}^{2}$. Construct the instance $(X-S_{V_{C}},r)$ for PCA.
@@ 2 Preliminaries
We use $[n]$ to denote the set ${1,2,\cdots,n}$ for a positive integer $n$. The distance of vector $x$ from subspace $V$ denoted by $\mathtt{dist}(x,V)$, is defined as $\inf_{v\in V}||x-v||$. While $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm of a matrix, which is the sum of the absolute values of its entries. Let $X\subset\mathbb{R}^{d}$ be a set of size $n$, we define $\mu(X)=\frac{1}{n}\sum_{x\in X}x$ as the mean of the vectors in $X$. In this paper, we consider the PCA with Outliers problem, which takes an instance $(X,r,k)$ as input, where $X$ is a subset of $\mathbb{R}^{d}$ of size $n$, and $r,k\in\mathbb{N}$. The goal is to find a $r$-dimensional subspace $L$ that captures the most important information contained in the data points of $X$ while ignoring up to $k$ outliers. These outliers, which are rows of $X$, will be denoted by a (sparse) matrix $S$ of the same size, where only the rows corresponding to the inliers are replaced with zeros. Throughout the paper, we will use matrix form and linear subspace spanned by rows of that matrix, interchangeably.
2.1 Probability Measure on Rank $r$ Subspaces
In this subsection, we establish a probability measure on the set of all rank $r$ subspaces of an $d$-dimensional real vector space. This set is formally represented by the Grassmannian manifold [^1] See [21] for more details on Grassmannian manifolds and their uses. $\text{Gr}(r,d)$, a fundamental object in differential geometry and algebraic topology.
The Grassmannian $\text{Gr}(r,d)$ is the space of all $r$-dimensional subspaces of $\mathbb{R}^{d}$. This manifold is both compact and connected, with its dimension given by:
$\dim\text{Gr}(r,d)=r(d-r).$
The Grassmannian is endowed with a natural topology and a differential structure, making it a Riemannian manifold. This structure allows for the application of various geometric and analytic techniques to study its properties.
To define a measure on $\text{Gr}(r,d)$, we utilize the Haar measure [^2] See [17] for definition and properties of the Haar measure.. The Haar measure is a unique, translation-invariant measure on the manifold, invariant under the action of the orthogonal group $O(d)$. This invariance is crucial as it ensures that the measure does not change under orthogonal transformations, preserving the intrinsic geometry of the manifold. The Haar measure is finite due to the compactness of the Grassmannian manifold, making it suitable for normalization.
To transform the Haar measure into a probability measure, we normalize it such that the total measure of the entire Grassmannian is 1. Let $\mu_{\text{Haar}}$ denote the Haar measure on $\text{Gr}(r,d)$. The normalized probability measure $\mu$ is then defined by:
$\mu(A)=\frac{\mu_{\text{Haar}}(A)}{\mu_{\text{Haar}}(\text{Gr}(r,d))},$
for any measurable subset $A\subset\text{Gr}(r,d)$. This normalization ensures that $\mu$ is a probability measure, i.e., $\mu(\text{Gr}(r,d))=1$.
Definition 1.
Formally, we define the probability measure $\mu$ on the Grassmannian manifold $\text{Gr}(r,d)$ as follows:
$\mu:\mathcal{B}(\text{Gr}(r,d))\to[0,1],$
where $\mathcal{B}(\text{Gr}(r,d))$ denotes the Borel $\sigma$-algebra [^3] For a comprehensive understanding of Borel $\sigma$-algebra, readers may look into the relevant sections in [25]. on $\text{Gr}(r,d)$, and
$\mu(A)=\frac{\mu_{\text{Haar}}(A)}{\mu_{\text{Haar}}(\text{Gr}(r,d))}.$
This construction ensures that $\mu$ is a well-defined probability measure, allowing for a normalized way to measure subsets of the Grassmannian manifold. This enables probabilistic analysis on the set of all rank $r$ subspaces of $\mathbb{R}^{d}$, facilitating the study of random subspaces.
By using this probability measure, we can rigorously analyze the likelihood of certain subspaces containing desired properties, such as minimizing projection distances for PCA. This approach is fundamental in developing algorithms that leverage random sampling of subspaces, providing a theoretical foundation for their effectiveness and efficiency.
Lemma 1.
The volume of the entire Grassmannian, denoted as $\text{Vol}(\text{Gr}(r,d))$, is given by:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
Proof.
The volume of the Grassmannian $\text{Gr}(r,d)$ is derived from the volumes of the associated manifold and orthogonal groups. The Grassmannian $\text{Gr}(r,d)$ can be seen as the quotient of the orthogonal group $O(d)$ by the product of the orthogonal groups $O(r)$ and $O(d-r)$. Therefore, its volume can be expressed in terms of these groups’ volumes.
In case of the orthogonal group $O(d)$, the volume is given by:
$\text{Vol}(O(d))=\frac{2^{d(d-1)/2}\pi^{d(d+1)/4}}{\prod_{j=1}^{d}\Gamma(j/2)}.$
In case of the manifold $V_{r}(\mathbb{R}^{d})$, the space of orthonormal vectors in $\mathbb{R}^{d}$, the volume is given by:
$\text{Vol}(V_{r}(\mathbb{R}^{d}))=\frac{\pi^{r(d-r)/2}}{\prod_{i=1}^{r}\Gamma((d-i+1)/2)}.$
The Grassmannian can be viewed as the quotient of the manifold by the orthogonal group $O(r)$:
$\text{Gr}(r,d)=\frac{V_{r}(\mathbb{R}^{d})}{O(r)}.$
Hence, the volume of $\text{Gr}(r,d)$ is:
$\text{Vol}(\text{Gr}(r,d))=\frac{\text{Vol}(V_{r}(\mathbb{R}^{d}))}{\text{Vol}(O(r))}.$
Substituting the volumes from the orthogonal group and manifold, we get:
$\text{Vol}(\text{Gr}(r,d))=\frac{\frac{\pi^{r(d-r)/2}}{\prod_{i=1}^{r}\Gamma((d-i+1)/2)}}{\frac{2^{r(r-1)/2}\pi^{r(r+1)/4}}{\prod_{j=1}^{r}\Gamma(j/2)}}.$
Simplifying this expression, we obtain:
$\text{Vol}(\text{Gr}(r,d))=\frac{\pi^{r(d-r)/2}\prod_{j=1}^{r}\Gamma(j/2)}{2^{r(r-1)/2}\pi^{r(r+1)/4}\prod_{i=1}^{r}\Gamma((d-i+1)/2)}.$
Further simplification leads to:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
Thus, we have derived the volume of the Grassmannian manifold $\text{Gr}(r,d)$. ∎
This measure offers a precise method for conducting probabilistic analyses of subspaces, which is essential for the effectiveness of algorithms that utilize random subspace sampling.
2.2 Voronoi diagrams
@@
Consider a space $X$ (e.g. $\mathbb{R}^{d}$ or a subset of it), a set of objects $S$ (e.g. a subset of $X$), and a distance function $d:X\times S\to\mathbb{R}$ on this space. A Voronoi diagram is a partitioning of $X$ into subsets each closest to the given set of objects $S$. Each element of the set $S$ such as $s_{i}$ has a corresponding region, called a Voronoi cell, consisting of all the elements of the space $X$ closer to that element $s_{i}$ than to any other element $s_{j\neq i}$ in the set. Voronoi diagrams have many applications in various fields [3, 23]. An example of a Voronoi diagram with a data set of size 15 is depicted in Figure 1.
@@
(Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Voronoi diagram is defined as the partitioning $V\subset\mathcal{P}(\mathbb{R}^{d})$ of $\mathbb{R}^{d}$ such that for a vector $s\in S$:
@@
Let $S$ be a set of $n$ subspaces of $\mathcal{L}(\mathbb{R}^{d})$ of rank $k$ and $X$ be the set of all rank $k$ subspaces of $\mathcal{L}(\mathbb{R}^{d})$. If we define distance $Gr_{k}:X\times S\to\mathbb{R}$ as the usual distance function of the Grassmannian, we can define the partitioning $V\subset\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ of $\mathcal{L}(\mathbb{R}^{d})$ such that for a subspaces $Y\in S$:
$\forall U,W\in\mathcal{L}(\mathbb{R}^{d})\quad U\sim W\iff\arg\min_{Y\in S}Gr_{k}(U,Y)=\arg\min_{Y\in S}Gr_{k}(W,Y)$
Although a normal Voronoi cell is defined as the set of elements closest to a single object in S, a $\mathtt{k}$th-degree Voronoi cell is defined as the set of elements having a particular tuple of $k$ objects in $S$ as their nearest neighbors. Higher-degree Voronoi diagrams also subdivide space.
Higher-degree Voronoi diagrams can be generated recursively. To generate the $\mathtt{k}$th-degree Voronoi diagram from set $S$, start with the $\mathtt{(k-1)}$th-degree diagram and replace each cell generated by $T=\{s_{1},s_{2},\cdots,s_{k-1}\}$ with a Voronoi diagram generated on the set $S\setminus T$.
@@
(Higher Degree Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $\mathtt{k}$th-degree Voronoi diagram as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathbb{R}^{d})$ such that:
$V(U)=\bigg{\{}\>\>x\>\>\bigg{|}\>\>\max_{u\in U}d(x,u)\leq\min_{t\in\mathcal{S}\setminus U}d(x,t)\>\>\land\>\>d(x,u_{1})\leq d(x,u_{2})\leq\cdots\leq d(x,u_{k})\bigg{\}}$
where $\mathcal{P}^{k}(A)$ is the set of all $\mathtt{k}$-tuples with elements from $A$ and $d(x,y)$ is the Euclidean distance function.
@@
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $\mathtt{k}$th-degree Voronoi diagram of hyperspaces with respect to vectors $\mathcal{S}$ as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ such that:
$V(U)=\bigg{\{}\>\>W\>\>\bigg{|}\>\>\max_{u\in U}d(W,u)\leq\min_{t\in\mathcal{S}\setminus U}d(W,t)\>\>\land\>\>d(W,u_{1})\leq d(W,u_{2})\leq\cdots\leq d(W,u_{k})\bigg{\}}$
where $\mathcal{P}^{k}(A)$ is the set of all $\mathtt{k}$-tuples with elements from $A$ and $d(x,y)$ is the Euclidean distance function.
@@
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Construction of the $\mathtt{k}$th-degree Voronoi diagram of hyperspaces with respect to vectors $\mathcal{S}$ as defined in definition 5 is computable in time $n^{O(d)}$.
@@
We are going to formulate the boundary of each cell of $\mathtt{(n-1)}$th-degree diagram with a set of linear equations which is going to result in $n^{2}$ hyperplanes. using proposition 4, we are able to compute the expected diagram in $n^{O(d)}$. $\mathtt{k}$th-degree Voronoi diagram is simply constructible by omitting the last $n-k$ indices of each cell.
For every two elements $s_{i}$ and $s_{j}$ from $\mathcal{S}$, Let $P_{ij}$ be a function on the set of all subspaces of the vector space $\mathcal{L}(\mathbb{R}^{d})$ to real numbers $\mathbb{R}$ which measures the difference in the distance of $s_{i}$ and $s_{j}$ to a given subspace $W$. So, if the projection operator on $W$ is $P_{W}$ then we can define this as follows:
$P_{ij}(W)=\|P_{W}\cdot s_{i}\|-\|P_{W}\cdot s_{j}\|$
By sorting the values of the $P_{ij}$’s for a specific $W$, we can determine the corresponding cell for this subspace. To elaborate further, since our focus lies only on the sequence of vector distances from $\mathcal{S}$ to $W$, the signs of the equations $P_{i,j}$ suffice. Consequently, a Voronoi cell can be represented by a sign condition based on these equations. Thus, by the continuity of the equations $P_{i,j}$ within the specified domain, we can deduce that a subspace $W$ lies on the boundary of a Voronoi cell only when there exist $i$ and $j$ such that $P_{i,j}(W)=0$. As we have $n^{2}$ equations with a unique rank $d-1$ solution, we have $n^{2}$ hyperspaces and their arrangments construct the expected diagram. This results in a $(n^{2})^{d}=n^{O(d)}$ time algorithm. ∎

### v3

@@ Abstract
In this paper, we introduce new algorithms for Principal Component Analysis (PCA) with outliers. Utilizing techniques from computational geometry, specifically higher-degree Voronoi diagrams, we navigate to the optimal subspace for PCA even in the presence of outliers. This approach achieves an optimal solution with time complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$. Additionally, we present a randomized algorithm with complexity $O(n^{r}\log(1/\delta)/C(d,r,\alpha))\text{poly}(n,d)$. Our approach leverages properties of high-dimensional spaces and the separation condition of outliers to efficiently recover the optimal subspace. Our results demonstrate that higher-degree Voronoi diagrams, combined with probabilistic subspace selection techniques, provide an effective and scalable solution for PCA with outliers.
@@ 1 Introduction
[…]
Here $\text{A}||_{F}^{2}=\sum_{i,j}{a_{i,j}^{2}}$ is the square of the Frobenius norm of matrix A.
@@
In the presence of outliers, the classical PCA algorithm may not work very well, as they can significantly affect the results. The PCA with Outliers problem attempts to improve the results of PCA by removing some data points that are considered outliers. The goal is to recover a low-rank matrix with sparse errors. Here, the assumption is that the data points lie near an $r$-dimensional subspace, and the remaining points are distributed randomly, possibly as a result of distorted observations. This problem can be formulated as follows: given a matrix $\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R})$, where $n$ is the number of data points and $d$ is the dimensionality of the data, and a parameter $r$ representing the rank of the low-rank matrix, the goal is to find a low-rank matrix $L$ and a sparse error matrix $S$ such that $\text{X}=S+L$, where $\text{rank}(L)\leq r$ and $S||_{\#}\leq k$, where $k$ is a parameter representing the number of outliers to be removed, and $||\cdot||_{\#}$ denotes the number of non-zero rows in a matrix, as shown in (2).
[…]
@@
Our Results The contribution of this paper is as follows.
First, we propose a new algorithm for PCA with Outliers with time complexity $n^{\mathcal{O}(d)}\text{poly}(n,d)$, which fills the gap identified in [27] as this is the optimal bound. To achieve this, we leverage higher-degree Voronoi diagrams, which are defined in the next section.
Second, we develop a randomized algorithm with time complexity $O\Bigl(\frac{n^{r}\log(1/\delta)}{C(d,r,\alpha)}\text{poly}(n,d)\Bigr),$ where $C(d,r,\alpha)$ is an explicit constant depending on $d,r,$ and the separation parameter $\alpha$. Our randomized algorithm leverages the $\alpha$-gap condition to ensure that inliers and outliers are well-separated, making it possible to identify the optimal subspace with high probability. We say the instance satisfies the $\alpha$-gap condition when every outlier is at least $(1+\alpha)$ times farther from the optimal subspace $L^{\star}$ than every inlier (see Lemma 1). Specifically, we prove that the probability of success is at least $1-\delta$, where $\delta$ controls the failure rate over multiple independent trials. The algorithm works by iteratively sampling random subspaces and selecting the one that best fits the inliers. This approach offers a scalable and conceptually intuitive solution to PCA with Outliers, particularly in high-dimensional settings where traditional deterministic methods are computationally infeasible. Our results indicate that higher-degree Voronoi diagrams, combined with randomized subspace selection, yield an optimal and practical solution for outlier-robust PCA.
@@ 1.1 Related works
To solve this problem exactly, one can trivially check all the possible $n\choose k$ ways to choose $k$ vectors as outliers and then perform PCA on the rest nkn-k vectors and return the best result among all this as the solution of PCA with Outliers. However, not all these $n\choose k$ configurations are possible when you consider the geometric properties of the given vectors and possible subspaces.
@@
Where $\|\cdot\|_{*}$ denotes the nuclear norm of a matrix, which is the sum of its singular values, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm of a matrix, which is the sum of the absolute values of its entries. The parameter $\lambda$ controls the trade-off between the low-rank and sparse components of the data matrix.
Also the PCA with Outliers problem can be solved using a variety of convex optimization techniques, such as the Alternating Direction Method of Multipliers or the Augmented Lagrangian Method. The paper [32] proposes the Outlier Pursuit method to solve this convex problem. This algorithm is based on the observation that the sparse matrix $S$ can be decomposed into a sum of sparse vectors, each of which corresponds to an outlier. It iteratively identifies and removes the outliers from the data matrix, and then computes the low-rank component of the remaining data using PCA. The outliers are then added back to the low-rank component to obtain the final estimate of the data matrix. The mentioned algorithm requires the sparsity level to be known a priori, which can be a limitation in practice. Additionally, the algorithm assumes that the noise in the data matrix is Gaussian, which may not be true in practice.
To overcome the issues of the previous methods, [11] proposed the Manipulator Pursuit algorithm, which is more robust and flexible. This method can handle wider range of outlier models and noise distributions. It does not require the sparsity level to be known a priori as it uses data-driven approach to estimate the sparsity level and a nuclear norm regularization term to encourage low-rank solutions.
In [27], the authors analyzed the computational complexity of the PCA with Outliers problem. They refined the lower bound on the time complexity of the problem by showing that it is $\Omega(n^{d})$. Finally it is concluded that the lower bound is tight up to a polynomial factor, and also an algorithm for solving this problem in $n^{\mathcal{O}(d^{2})}{\text{poly}(n,d)}$ is proposed using tools from real algebraic geometry. To be explicit, their result is summarized in the following proposition.
@@
([6], Theorem 13.22) Consider $V$ as an algebraic set in $\mathbb{R}^{d}$ of real dimension $d^{\prime}$, defined by $Q(X_{1},\ldots,X_{d})=0$, where $Q$ is a polynomial in $\mathbb{R}[X_{1},\ldots,X_{d}]$ of degree at most $b$. Let $\mathcal{P}\subset\mathbb{R}[X_{1},\ldots,X_{d}]$ be a finite set of $s$ polynomials, each $P\in\mathcal{P}$ also having a degree at most $b$. Define $D$ as the ring generated by the coefficients of $Q$ and the polynomials in $\mathcal{P}$. Given $Q$, $d^{\prime}$, and $\mathcal{P}$, there exists an algorithm that computes a set of points such that each non-empty cell of $V$ over $\mathcal{P}$ contains at least one point from the set. This algorithm requires at most $s^{d^{\prime}}b^{O(d)}$ arithmetic operations in $D$.
@@
    For each $V_{C}$, compute the optimal $S_{V_{C}}$ by selecting the $k$ rows of $X$ with the highest value of $V_{C}x^{T}_{i}||_{F}^{2}$. Construct the instance $(X-S_{V_{C}},r)$ for PCA.
@@ 2 Notation
In this paper, we consider the PCA with Outliers problem, which takes an instance $(X,r,k)$ as input, where $X$ is a subset of $\mathbb{R}^{d}$ of size $n$, and $r,k\in\mathbb{N}$. The goal is to find a $r$-dimensional subspace $L$ that captures the most important information contained in the data points of $X$ while ignoring up to $k$ outliers. These outliers, which are rows of $X$, will be denoted by a (sparse) matrix $S$ of the same size, where only the rows corresponding to the inliers are replaced with zeros. Throughout the paper, we will use matrix form and linear subspace spanned by rows of that matrix, interchangeably. Also, for $x\in\mathbb{R}^{d}$ and $\epsilon>0$, we denote by $B_{\epsilon}^{d}(x)\;=\;\{\,y\in\mathbb{R}^{d}:\|y-x\|_{2}\leq\epsilon\}$ the closed Euclidean ball of radius $\epsilon$ centered at $x$ in $d$-dimensional space.
3 Voronoi diagrams
@@
Consider a space $X$ (e.g. $\mathbb{R}^{d}$ or a subset of it), a set of objects $S$ (e.g. a subset of $X$), and a distance function $d:X\times S\to\mathbb{R}$ on this space. A Voronoi diagram partitions $X$ into regions, each corresponding to the set of points closest to an element of $S$. Each element of the set $S$ such as $s_{i}$ has a corresponding region, called a Voronoi cell, consisting of all the elements of the space $X$ closer to that element $s_{i}$ than to any other element $s_{j\neq i}$ in the set. Voronoi diagrams have many applications in various fields [4, 25]. An example of a Voronoi diagram with a data set of size 15 is depicted in Figure 1.
@@
(Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Voronoi diagram is defined as the partitioning $V\subset\mathcal{P}(\mathbb{R}^{d})$ of $\mathbb{R}^{d}$ such that for a vector $s\in\mathcal{S}$:
@@
Let $\mathcal{S}$ be a set of $n$ vectors in $\mathbb{R}^{d}$. We can define the partitioning $V\subset\mathcal{P}(\mathcal{L}_{r}(\mathbb{R}^{d}))$ of $\mathcal{L}_{r}(\mathbb{R}^{d})$ such that for a vector $s\in\mathcal{S}$:
$\forall U,V\in\mathcal{L}_{r}(\mathbb{R}^{d})\quad U\sim V\iff\arg\min_{s\in\mathcal{S}}\mathtt{dist}(x,U)=\arg\min_{s\in\mathcal{S}}\mathtt{dist}(x,V)$
Although a normal Voronoi cell is defined as the set of elements closest to a single object in $S$, a $\mathtt{k}$-degree Voronoi cell is defined as the set of elements having a particular tuple of $k$ objects in $S$ as their nearest neighbors. They also subdivide space.
Moreover, higher-degree Voronoi diagrams can be generated recursively. To generate the $\mathtt{k}$-degree Voronoi diagram from set $S$, start with the $\mathtt{(k-1)}$th-degree diagram and replace each cell generated by $T=\{s_{1},s_{2},\cdots,s_{k-1}\}$ with a Voronoi diagram generated on the set $S\setminus T$.
@@
(Higher Degree Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $k$-degree Voronoi diagram as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathbb{R}^{d})$ such that:
$V(U)=\bigg\{\>\>x\>\>\bigg|\>\>\max_{u\in U}\mathtt{dist}(x,u)\leq\min_{t\in\mathcal{S}\setminus U}\|x-u\|\>\>\land\>\>\|x-u_{1}\|\leq\|x-u_{2}\|\leq\cdots\leq\|x-u_{k}\|\bigg\}$
where $\mathcal{P}^{k}(A)$ is the set of all $k$-tuples with elements from $A$ and $U=(u_{1},u_{2},\ldots,u_{k})$ is an ordered $k$-tuple of points from $\mathcal{S}$ (i.e. $U\in\mathcal{P}^{k}(\mathcal{S})$).
@@
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $k$-degree Voronoi diagram of $r$-dimensional subspaces with respect to vectors $\mathcal{S}$ as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ such that:
$\begin{split}V(U)=\bigg\{\>\>W\>\>\bigg|&\>\>\max_{u\in U}\mathtt{dist}(u,W)\leq\min_{t\in\mathcal{S}\setminus U}\mathtt{dist}(t,W)\\ &\>\>\land\>\>\mathtt{dist}(u_{1},W)\leq\mathtt{dist}(u_{2},W)\leq\cdots\leq\mathtt{dist}(u_{k},W)\bigg\}\end{split}$
where $\mathcal{P}^{k}(A)$ is the set of all $k$-tuples with elements from $A$ and $\mathtt{dist}(x,y)$ is the Euclidean distance function and $\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ is the set of all linear subspaces of $\mathbb{R}^{d}$.
As can be seen, in the context of the PCA with Outliers problem, the order of the data is not important relative to the desired subspace. However, our challenge was that we could not directly represent the orderless partitioning of $\mathbb{R}^{n}$ for subspaces in a suitable arrangement of hyperplanes to apply the proposition 4, and therefore we presented a formulation of the higher degree diagrams on the subspace.
@@
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Construction of the k{k}-degree Voronoi diagram of hyperspaces with respect to vectors $\mathcal{S}$ as defined in definition 4 is computable in time $n^{O(d)}$.
@@
We are going to formulate the boundary of each cell of (n1){(n-1)}-degree diagram with a set of linear equations which is going to result in $n^{2}$ hyperplanes. using proposition 4, we are able to compute the expected diagram in $n^{O(d)}$. k{k}-degree Voronoi diagram is simply constructible by omitting the last nkn-k indices of each cell.
For every two elements $s_{i}$ and $s_{j}$ from $\mathcal{S}$, Let $P_{i,j}$ be a function on the set of all subspaces of the vector space $\mathcal{L}(\mathbb{R}^{d})$ to real numbers $\mathbb{R}$ which measures the difference in the distance of $s_{i}$ and $s_{j}$ to a given subspace $W$. So, if the projection operator on $W$ is $P_{W}$ then we can define this as follows:
$P_{i,j}(W)=\|P_{W}\cdot s_{i}\|-\|P_{W}\cdot s_{j}\|$
By sorting the values of the $P_{i,j}$’s for a specific $W$, we can determine the corresponding cell for this subspace. To elaborate further, since our focus lies only on the sequence of vector distances from $\mathcal{S}$ to $W$, the signs of the equations $P_{i,j}$ suffice. Consequently, a Voronoi cell can be represented by a sign condition based on these equations. Thus, by the continuity of the equations $P_{i,j}$ within the specified domain, we can deduce that a subspace $W$ lies on the boundary of a Voronoi cell only when there exist $i$ and $j$ such that $P_{i,j}(W)=0$. As we have $n^{2}$ equations with a unique rank d1d-1 solution, we have $n^{2}$ hyperspaces and their arrangments construct the expected diagram. This results in a $(n^{2})^{d}=n^{O(d)}$ time algorithm. ∎